what is The pH of buffer prepared by adding 10ml of 0.10M acetic acid to 20ml of
e. 0.10MCH3COONa. (Ka=1.75x10-5) ?
1
Expert's answer
2020-09-21T06:41:17-0400

"pH=pK_a+\log\frac{[salt]}{[acid]}\\pH=4.756+\log\frac{20\times0.1}{10\times0.1}=4.756+\log2=5.056"